OpenStudy (anonymous):

For the function rule y=-2x-3, provide the x-y coordinates of the y-intercept and x-intercept.

jimthompson5910 (jim_thompson5910):

x-intercept: Plug in y = 0 and solve for x y=-2x-3 0=-2x-3 ...Tell me what you get ------------------------------------------------------- y-intercept Plug in x = 0 and evaluate to find y y=-2x-3 y=-2(0)-3 ...and tell me what you get here

OpenStudy (anonymous):

i dont understand it

jimthompson5910 (jim_thompson5910):

Do you know how to solve 0=-2x-3 for x?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

Alright, I'll walk you through it 0=-2x-3 0+3=-2x-3+3 ... Add 3 to both sides 3 = -2x ... Notice how -3+3 = 0 (and goes away) 3/(-2) = (-2x)/(-2) ... Divide both sides by -2 -3/2 = x ... The -2 terms on the right side divide and cancel (and go away) x = -3/2 ... Flip the equation So when y = 0, x is x = -3/2 So the x-intercept is \(\Large \left(-\frac{3}{2},0\right)\) Does that make sense?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

ok great

jimthompson5910 (jim_thompson5910):

Can you tell me what y=-2(0)-3 becomes ?

OpenStudy (anonymous):

-1?

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

y=-2(0)-3 y=0-3 .... negative 2 times zero is zero y = -3 ... subtract So y=-2(0)-3 becomes y = -3

jimthompson5910 (jim_thompson5910):

This means that when x = 0, y is y = -3

jimthompson5910 (jim_thompson5910):

So the y-intercept is \(\Large \left(0,-3\right)\)

OpenStudy (anonymous):

OHHH ok now i see

jimthompson5910 (jim_thompson5910):

that's great

OpenStudy (anonymous):

Now what about the x intercept?

jimthompson5910 (jim_thompson5910):

Look a few posts above and you'll see the work, but the x-intercept again is \(\Large \left(-\frac{3}{2},0\right)\)

OpenStudy (anonymous):

ok sorry i looked past it

jimthompson5910 (jim_thompson5910):

that's ok

OpenStudy (anonymous):

Thanks

jimthompson5910 (jim_thompson5910):

you're welcome